# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 3.1.5 Tesauro (1994)
Backgammon was far beyond tabular methods, yet TD-Gammon trained a feedforward neural network to estimate the probability of winning from any board position. A hidden layer of 80 sigmoid units fed a single sigmoid output:
@@
where the eligibility trace $\mathbf{e}_{t}=\sum_{k=1}^{t}\lambda^{t-k}\nabla_{\boldsymbol{\theta}}\hat{V}(s_{k})$ accumulates exponentially decayed gradients of past predictions. [^16] In the neural network case, the eligibility trace $\mathbf{e}_{t}\in\mathbb{R}^{|\boldsymbol{\theta}|}$ is a vector accumulating exponentially-weighted gradients, extending the scalar state-based trace to parameter space. At game’s end, $\hat{V}(s_{t+1})$ is replaced by the outcome $z\in\{0,1\}$.
Self-play proceeds as follows. A single neural network $\hat{V}$ serves as the evaluation function for both players. At each turn, the current player rolls the dice, which determines the set of legal moves. For each legal move, the network evaluates the resulting board position $\hat{V}(s^{\prime})$. The player selects the move that maximizes $\hat{V}(s^{\prime})$ from its own perspective. The opponent then does the same using the identical network. After each move from position $s_{t}$ to $s_{t+1}$, the weights are updated by the TD($\lambda$) rule using the error $\hat{V}(s_{t+1})-\hat{V}(s_{t})$, which pushes earlier predictions toward later ones. When the game ends, the terminal signal $z\in\{0,1\}$ replaces $\hat{V}(s_{t+1})$, grounding the entire chain of predictions in the actual outcome. As the network improves, it generates stronger play on both sides, which produces harder training games, which in turn drives further improvement. The dice rolls ensure that the network encounters a wide variety of board positions without requiring an explicit exploration mechanism [^17] TD-Gammon 1.0, trained on 300,000 self-play games, played at a strong intermediate level. Version 1.0 trained on 300,000 self-play games with 80 hidden units and no lookahead search; version 2.1 trained on 1,500,000 games with 2-ply search, reducing its average loss rate against world-class human players from 0.25-0.25 to 0.02-0.02 points per game. Version 2.1 achieved near-parity with Bill Robertie, a former world champion. The program discovered novel positional strategies subsequently adopted by the human backgammon community..
@@ 3.1.6 SARSA (1994)
Q-learning learns the optimal action-value function regardless of the policy generating experience. This off-policy property is useful but introduces complications when combined with function approximation. Rummery and Niranjan (1994) introduced SARSA as an on-policy [^18] On-policy methods learn about the policy currently generating data, in contrast to off-policy methods (see footnote on Q-learning above) which learn about a different target policy. alternative that learns the value of the policy actually being followed. The name derives from the quintuple $(s_{t},a_{t},r_{t+1},s_{t+1},a_{t+1})$ used in each update:
@@ 3.1.8 Actor-Critic Methods (2000)
The idea of maintaining both a policy (actor) and a value function (critic) dates to Barto et al. (1983), who used a two-component system to solve the pole-balancing task. The actor selected actions; the critic predicted future reinforcement and generated training signals for the actor. This architecture addressed a limitation of pure policy gradient methods. REINFORCE estimates gradients from Monte Carlo returns, which are unbiased but have high variance because returns depend on the full trajectory of stochastic transitions and rewards. TD methods use bootstrapped targets $r+\gamma V(s^{\prime})$, which have lower variance but introduce bias when $V$ is approximate.
Actor-critic methods combine both ideas. The critic learns a value function $V(s)$ by TD updates:
@@
The actor updates the policy using the TD error as a sample of the advantage:
@@ 3.2.1 Deep Q-Networks (2015)
The two decades after TD-Gammon saw limited success in applying neural networks to reinforcement learning. Attempts to replicate Tesauro’s approach in chess and Go did not achieve comparable results, though progress continued in robotics and other domains. Mnih et al. (2015) changed this by training a single convolutional neural network [^22] A convolutional neural network applies learned spatial filters to detect local patterns in grid-structured data, commonly used for image inputs. to play 49 Atari 2600 games directly from pixel inputs ($210\times 160\times 3$) and a scalar score, using no game-specific features.
The architecture processed four consecutive frames through three convolutional layers and a fully connected layer. The network $Q(s,a;\theta)$ was trained to minimize the squared temporal difference error:
@@
Two innovations stabilized learning. Experience replay (Lin, 1992) stored transitions $(s,a,r,s^{\prime})$ in a buffer $\mathcal{D}$ holding $10^{6}$ transitions and sampled uniformly for training. This breaks the temporal correlation between consecutive updates that would otherwise cause the network to overfit to recent experience. A target network used a frozen copy of parameters $\theta^{-}$, [^23] The notation $\theta^{-}$ denotes target network parameters, a lagged copy of $\theta$ updated periodically. This stabilizes learning by keeping the regression target fixed during gradient steps. updated to match $\theta$ only every C=10, [^0] C=10{,}000 steps, so the regression target does not shift with each gradient step.
DQN exceeded human-level performance on 29 of 49 games using a single architecture and hyperparameters. The same network that learned Breakout (where the agent must track a ball and position a paddle) also learned Space Invaders (where the agent must track multiple enemies and time shots). Games requiring long-horizon planning or sparse rewards, such as Montezuma’s Revenge (where the agent must navigate rooms and collect keys before receiving any reward), remained difficult.
@@ 3.2.2 TRPO and PPO (2015, 2017)
Policy gradient methods suffer from a practical instability. A single large gradient step can move the policy into a region of parameter space where performance collapses, and recovery may be slow or impossible. The objective $J(\theta)$ is not well approximated by its linear Taylor expansion when the step size is large relative to the curvature of the policy space.
@@
where $L$ is a surrogate objective based on the advantage function $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$. [^24] The Kullback-Leibler divergence $D_{\mathrm{KL}}(p\|q)=\sum_{x}p(x)\log(p(x)/q(x))$ measures statistical distance between distributions. The bar denotes expectation over states: $\bar{D}_{\mathrm{KL}}=\mathbb{E}_{s}[D_{\mathrm{KL}}(\pi_{\theta_{\text{old}}}(\cdot|s)\|\pi_{\theta}(\cdot|s))]$. The KL constraint prevents the policy from moving so far in one update that performance collapses and recovery becomes impossible.
@@ 3.2.4 AlphaGo Zero (2017)
AlphaGo Zero uses a single neural network $f_{\theta}(s)=(\mathbf{p},v)$ that takes a board position $s$ and outputs both a policy vector $\mathbf{p}$ over legal moves and a scalar value $v$ estimating the probability of winning. [^25] A convolutional neural network applies small learned filters that slide across a grid input, detecting local patterns regardless of position on the board. The same filter that recognizes a capturing pattern in one corner will recognize it in another. The input representation consists of 17 binary planes on the $19\times 19$ board, encoding the raw game state without any hand-crafted features: eight planes for Black’s stone positions over the last eight moves, eight for White’s, and one indicating which color plays next. [^26] The history planes allow the network to detect ko situations (positions that cannot be immediately repeated) and infer the trajectory of play without any explicit feature engineering. The board’s fixed grid structure makes convolutional networks effective, since spatial patterns learned in one region transfer to others. The architecture uses residual blocks, [^27] A residual block computes $\mathbf{x}+g(\mathbf{x})$ rather than just $g(\mathbf{x})$, where $g$ is a learned transformation. The skip connection allows gradients to flow through very deep networks without vanishing. AlphaGo Zero’s 40-block architecture has 79 parameterized layers. which allow training of very deep networks.
During play, each move is selected by running MCTS, which conducts 1,600 simulated games from the current position to estimate move quality. Each simulation proceeds in four phases, illustrated in Figure 1. In the selection phase, the algorithm starts from the current position and traverses the partially built search tree by choosing at each node the action that maximizes $Q(s,a)+c_{\text{puct}}\cdot P(s,a)\cdot\sqrt{\sum_{b}N(s,b)}\,/\,(1+N(s,a))$, where $Q(s,a)$ is the current average value of action $a$, $P(s,a)$ is the prior probability from the neural network, and $N(s,a)$ is the visit count. [^28] The constant $c_{\text{puct}}$ controls exploration. Actions visited often have well-estimated $Q$ values but a shrinking exploration bonus; rarely visited actions have uncertain values but a large bonus. This is a continuous analogue of the upper confidence bound (UCB) strategy from bandit theory. In the expansion and evaluation phase, when the traversal reaches a position not yet in the tree, the neural network evaluates it in a single forward pass, [^29] A forward pass is one evaluation of the neural network on an input, producing outputs without any parameter updates. producing a policy vector $\mathbf{p}$ and a value estimate $v$; the policy initializes prior probabilities $P(s^{\prime},a)=p_{a}$ for each child edge. In the backup phase, the value $v$ propagates back up the traversed path, incrementing each edge’s visit count $N(s,a)$ and updating its mean value $Q(s,a)$. After all 1,600 simulations, the algorithm selects the move with the highest visit count at the root.
@@ 4.1.1 Value Iteration as Picard Iteration
Consider the Bellman optimality operator $T$ acting on value functions:
@@ 4.1.2 Policy Iteration as Newton’s Method
Policy iteration takes a different approach. At the current value estimate $\tilde{V}$, define the greedy policy $\tilde{\pi}(s)=\mathop{\it argmax}_{a}\{r(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)\tilde{V}(s^{\prime})\}$. The policy evaluation step solves the linear fixed-point equation $V=T^{\tilde{\pi}}V$ exactly, where $T^{\tilde{\pi}}$ is the policy-specific Bellman operator:
@@
The consequence is finite termination. Since there are at most $|\mathcal{A}|^{|\mathcal{S}|}$ deterministic policies and each PI step strictly improves the value function (unless the current policy is already optimal), PI reaches the exact optimum in finitely many iterations. While VI requires $k=\log(100)/\log(1/\gamma)$ iterations to reduce error by a factor of 100 (Bertsekas, 1996), PI typically converges in 5–10 iterations regardless of $\gamma$. [^38] Ye (2011) proves PI is strongly polynomial with iteration count $O\!\left(\frac{|\mathcal{S}||\mathcal{A}|}{1-\gamma}\log\frac{|\mathcal{S}|}{1-\gamma}\right)$, resolving a long-standing conjecture. This bound is for fixed $\gamma$; Fearnley (2010) constructs examples requiring exponentially many iterations when $\gamma$ is allowed to vary with $|\mathcal{S}|$. [^39] For continuous-state problems discretized on a grid (the norm in economics), the finite-termination argument still applies to the discretized problem, but the number of grid points $n$ enters the bound. Santos and Rust (2004) establish a three-tier convergence result for PI applied to discretized dynamic programs: order $\approx 1.5$ globally for general interpolation schemes, quadratic convergence locally when the value function approximation is concave and piecewise linear, and superlinear convergence for general smooth interpolation. The formal error constants $C(h)$ in their quadratic bound degrade as the grid mesh $h\to 0$, but the iteration count is empirically independent of grid size. At $\gamma=0.90$, VI needs 44 iterations while PI needs only 5–8; at $\gamma=0.95$, VI requires 90 versus 5–8; at $\gamma=0.99$, VI needs 459 iterations while PI still converges in 5–10.
@@
Bertsekas (2022b) extends this interpretation to a broad class of dynamic programming problems. The Newton structure applies whenever the Bellman operator can be written as a pointwise maximum over linear operators: $T=\max_{\pi}T^{\pi}$. This includes not only infinite horizon problems with discounting but also optimal stopping problems (job search, option exercise), average cost optimization (inventory, queueing), and minimax formulations for adversarial settings. [^40] These problems appear under different names such as “stochastic shortest path” for optimal stopping, “average-cost MDP” for long-run average optimization, and “model predictive control” (or receding-horizon control) for finite-horizon replanning. The practical implication is that algorithms with policy-improvement structure (evaluate a policy exactly or approximately, then improve) inherit Newton-like convergence behavior, while pure value-iteration methods (apply $T$ directly) are limited to linear convergence. [^41] Blackwell (1965) proves that for discounted MDPs with finite state and action spaces, a stationary deterministic policy $\pi:\mathcal{S}\to\mathcal{A}$ exists that is optimal for all initial states simultaneously. This “uniform optimality” has three implications: (1) the search space reduces from history-dependent or stochastic policies to static maps, justifying neural networks that condition only on current state; (2) the optimal policy is independent of the initial distribution $d_{0}$, so changing where episodes start does not require retraining; (3) PI and VI are guaranteed to converge to the same globally optimal policy regardless of initialization.

### 15 Mar 2026 v2

@@ 4.1.5 Tesauro (1994)
Backgammon was far beyond tabular methods, yet TD-Gammon trained a feedforward neural network to estimate the probability of winning from any board position. A hidden layer [^19] A feedforward neural network stacks an input layer, one or more hidden layers, and an output layer; each unit in a hidden layer computes a weighted sum of its inputs and applies a nonlinear activation, here the logistic sigmoid $\sigma(x)=1/(1+e^{-x})$, which maps any real number to $(0,1)$ and is identical to the binary logit link function. of 80 sigmoid units fed a single sigmoid output.
@@
where $\alpha$ [^21] The learning rate $\alpha$ controls the step size of each parameter update. TD learning uses a semi-gradient step rather than true gradient descent, but $\alpha$ plays the same role: too large and updates overshoot; too small and convergence is slow. is the step size, and the eligibility trace $\mathbf{e}_{t}=\sum_{k=1}^{t}\lambda^{t-k}\nabla_{\boldsymbol{\theta}}\hat{V}(s_{k})$ accumulates exponentially decayed gradients of past predictions. [^22] In the neural network case, the eligibility trace $\mathbf{e}_{t}\in\mathbb{R}^{|\boldsymbol{\theta}|}$ is a vector accumulating exponentially-weighted gradients, extending the scalar state-based trace to parameter space. At game’s end, $\hat{V}(s_{t+1})$ is replaced by the outcome $z\in\{0,1\}$.
A single neural network $\hat{V}$ serves as the evaluation function for both players. At each turn, the current player selects the legal move maximizing $\hat{V}(s^{\prime})$ from its own perspective. As the network improves, it generates stronger play on both sides, producing harder training games that drive further improvement. The dice rolls ensure diverse board positions without requiring an explicit exploration mechanism. [^23] Version 2.1, trained on 1,500,000 games with 2-ply search, achieved near-parity with former world champion Bill Robertie and discovered novel positional strategies subsequently adopted by the human backgammon community.
@@ 4.1.6 SARSA (1994)
Q-learning learns the optimal action-value function regardless of the policy generating experience. This off-policy property is useful but introduces complications when combined with function approximation. Rummery and Niranjan (1994) introduced SARSA as an on-policy [^24] See Section 2 for the on-policy/off-policy distinction. alternative that learns the value of the policy actually being followed. The name derives from the quintuple $(s_{t},a_{t},r_{t+1},s_{t+1},a_{t+1})$ used in each update:
@@ 4.1.8 Actor-Critic Methods (2000)
The idea of maintaining both a policy (actor) and a value function (critic) dates to Barto et al. (1983), who used a two-component system to solve the pole-balancing task. Actor-critic methods address the high variance of REINFORCE by replacing Monte Carlo returns with bootstrapped TD targets as the learning signal. The critic learns a value function $V(s)$ by TD updates.
@@
The actor updates the policy using the TD error as a sample of the advantage.
@@ 4.1.9 Natural Policy Gradient (2001)
4.1.10 Fitted Value Iteration and Fitted Q-Iteration (2005)
Tabular Q-learning maintains a separate entry for every state-action pair. When $|\mathcal{S}|$ is large or the state space is continuous, as in most economic applications, this is infeasible. Fitted Q-Iteration (FQI) (Ernst et al., 2005) replaces the tabular update with a supervised regression step: given a batch of transitions, fit a function approximator to the Bellman targets.
Let $\mathcal{F}$ be a function class mapping $\mathcal{S}\times\mathcal{A}\to\mathbb{R}$. Initialize $Q_{0}\equiv 0$. At each iteration $k=0,1,\ldots,K-1$: (i) draw $N$ transitions $(s_{i},a_{i},r_{i},s_{i}^{\prime})$ from a generative model; (ii) construct regression targets $y_{i}^{(k)}=r_{i}+\gamma\max_{a^{\prime}}Q_{k}(s_{i}^{\prime},a^{\prime})$; (iii) set $Q_{k+1}\leftarrow\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{i=1}^{N}\bigl(f(s_{i},a_{i})-y_{i}^{(k)}\bigr)^{2}$. The output is the greedy policy $\pi_{K}(s)=\arg\max_{a}Q_{K}(s,a)$.
The regression step replaces the exact Bellman application with a projection onto $\mathcal{F}$: $Q_{k+1}=\Pi_{\mathcal{F}}\mathcal{T}Q_{k}$, where $\mathcal{T}$ is the Bellman optimality operator and $\Pi_{\mathcal{F}}$ is the $L^{2}$-projection under the sample distribution. Fitted Value Iteration (FVI) (Munos and Szepesvári, 2008) applies the same idea to the value function directly: $V_{k+1}=\Pi_{\mathcal{F}}\mathcal{T}^{*}V_{k}$, where $(\mathcal{T}^{*}V)(s)=\max_{a}\{r(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)V(s^{\prime})\}$.
With feature matrix $\Phi\in\mathbb{R}^{|\mathcal{S}|\times d}$ (rows $\phi(s)^{\top}$) and per-action weight vectors $\theta_{a}\in\mathbb{R}^{d}$, each FQI regression step for action $a$ reduces to the normal equations:
$\theta_{a}^{(k+1)}=\bigl(\Phi^{\top}\Phi\bigr)^{-1}\Phi^{\top}y_{a}^{(k)},$ (22)
where $y_{a}^{(k)}(s)=r(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)\max_{a^{\prime}}\phi(s^{\prime})^{\top}\theta_{a^{\prime}}^{(k)}$. Computation is $O(d^{2}|\mathcal{S}|+d^{3})$ per action per iteration. The FVI update takes the same form with a single weight vector $\theta_{V}$:
$\theta_{V}^{(k+1)}=\bigl(\Phi^{\top}\Phi\bigr)^{-1}\Phi^{\top}V_{\mathrm{target}}^{(k)},$ (23)
where $V_{\mathrm{target}}^{(k)}(s)=\max_{a}\bigl\{r(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)\phi(s^{\prime})^{\top}\theta_{V}^{(k)}\bigr\}$. The finite-sample error theory for these methods is developed in Section 5.2.5.
@@ 4.2.1 Deep Q-Networks (2015)
Mnih et al. (2015) trained a single convolutional neural network [^28] A convolutional neural network applies learned spatial filters to detect local patterns in grid-structured data, commonly used for image inputs. to play 49 Atari 2600 games directly from pixel inputs ($210\times 160\times 3$) and a scalar score, using no game-specific features.
The architecture processed four consecutive frames through three convolutional layers and a fully connected layer. [^29] A fully connected layer computes $y=Wx+b$ where every input unit is connected to every output unit with learned weights $W$ and biases $b$; it is the affine transformation familiar from linear regression, followed by a nonlinear activation. The network $Q(s,a;\theta)$ was trained to minimize the squared temporal difference error
@@
Two innovations stabilized learning. Experience replay (Lin, 1992) stored transitions $(s,a,r,s^{\prime})$ in a buffer $\mathcal{D}$ and sampled uniformly for training, breaking the temporal correlation between consecutive updates. A target network used a frozen copy of parameters $\theta^{-}$, updated periodically, [^30] The buffer held $10^{6}$ transitions; the target network $\theta^{-}$ was synchronized to $\theta$ every C=10, [^0] C=10{,}000 steps. so the regression target does not shift with each gradient step.
DQN exceeded human-level performance on 29 of 49 games using a single architecture and hyperparameters. [^31] Hyperparameters are design choices fixed before training begins, such as network depth, learning rate, and replay buffer size; they are not estimated by gradient descent and are analogous to tuning parameters in nonparametric estimation (bandwidth, penalty weights). Games requiring long-horizon planning or sparse rewards, such as Montezuma’s Revenge, remained difficult.
@@ 4.2.2 TRPO and PPO (2015, 2017)
Policy gradient methods suffer from a practical instability: a single large gradient step can move the policy into a region where performance collapses and recovery is slow.
@@
where $L$ is a surrogate objective based on the advantage function $A^{\pi}(s,a)=Q^{\pi}(s,a)-V^{\pi}(s)$. [^32] The Kullback-Leibler divergence $D_{\mathrm{KL}}(p\|q)=\sum_{x}p(x)\log(p(x)/q(x))$ measures statistical distance between distributions. The bar denotes expectation over states: $\bar{D}_{\mathrm{KL}}=\mathbb{E}_{s}[D_{\mathrm{KL}}(\pi_{\theta_{\text{old}}}(\cdot|s)\|\pi_{\theta}(\cdot|s))]$.
@@ 4.2.3 Soft Actor-Critic (2018)
Figure 1: Architecture comparison of the three fundamental algorithm families. (a) DQN maps states to Q-values for all actions, selecting the argmax. (b) REINFORCE maps states to a probability distribution over actions, then samples. (c) Actor-Critic maintains separate policy and value networks; the critic’s TD error $\delta_{t}$ provides a low-variance learning signal to the actor.
@@ 4.2.4 AlphaGo Zero (2017)
AlphaGo Zero uses a single convolutional neural network $f_{\theta}(s)=(\mathbf{p},v)$ that takes a board position $s$ and outputs both a policy vector $\mathbf{p}$ over legal moves and a scalar value $v$ estimating the probability of winning. The input representation consists of 17 binary planes on the $19\times 19$ board encoding the raw game state without hand-crafted features. [^33] Eight planes for Black’s stone positions over the last eight moves, eight for White’s, and one indicating which color plays next. The history planes allow the network to detect ko situations and infer the trajectory of play. The architecture uses residual blocks, [^34] A residual block computes $\mathbf{x}+g(\mathbf{x})$ rather than just $g(\mathbf{x})$, where $g$ is a learned transformation. The skip connection allows gradients to flow through very deep networks without vanishing. AlphaGo Zero’s 40-block architecture has 79 parameterized layers. which allow training of very deep networks.
During play, each move is selected by running MCTS, which conducts 1,600 simulated games from the current position to estimate move quality. Each simulation proceeds in four phases, illustrated in Figure 2. In the selection phase, the algorithm starts from the current position and traverses the partially built search tree by choosing at each node the action that maximizes $Q(s,a)+c_{\text{puct}}\cdot P(s,a)\cdot\sqrt{\sum_{b}N(s,b)}\,/\,(1+N(s,a))$, where $Q(s,a)$ is the current average value of action $a$, $P(s,a)$ is the prior probability from the neural network, and $N(s,a)$ is the visit count. [^35] The constant $c_{\text{puct}}$ controls exploration. Actions visited often have well-estimated $Q$ values but a shrinking exploration bonus; rarely visited actions have uncertain values but a large bonus. This is a continuous analogue of the upper confidence bound (UCB) strategy from bandit theory. In the expansion and evaluation phase, when the traversal reaches a position not yet in the tree, the neural network evaluates it in a single forward pass, producing a policy vector $\mathbf{p}$ and a value estimate $v$; the policy initializes prior probabilities $P(s^{\prime},a)=p_{a}$ for each child edge. In the backup phase, the value $v$ propagates back up the traversed path, incrementing each edge’s visit count $N(s,a)$ and updating its mean value $Q(s,a)$. After all 1,600 simulations, the algorithm selects the move with the highest visit count at the root.
@@ 5.1.1 Value Iteration as Picard Iteration
Consider the Bellman optimality operator $T$ acting on value functions.
@@ 5.1.2 Policy Iteration as Newton’s Method
Policy iteration takes a different approach. At the current value estimate $\tilde{V}$, define the greedy policy $\tilde{\pi}(s)=\mathop{\it argmax}_{a}\{r(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)\tilde{V}(s^{\prime})\}$. The policy evaluation step solves the linear fixed-point equation $V=T^{\tilde{\pi}}V$ exactly, where $T^{\tilde{\pi}}$ is the policy-specific Bellman operator
@@
Theorem 1 (Policy Improvement, Howard (1960)).
Let $\pi_{k}$ be the current policy with value $V^{\pi_{k}}$, and let $\pi_{k+1}$ be the greedy policy with respect to $V^{\pi_{k}}$:
$\pi_{k+1}(s)=\mathop{\it argmax}_{a\in\mathcal{A}}\left\{r(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)\,V^{\pi_{k}}(s^{\prime})\right\}.$ (32)
Then $V^{\pi_{k+1}}(s)\geq V^{\pi_{k}}(s)$ for all $s\in\mathcal{S}$, with strict inequality at some state unless $\pi_{k}$ is already optimal.
@@
The consequence is finite termination. Since there are at most $|\mathcal{A}|^{|\mathcal{S}|}$ deterministic policies and each PI step strictly improves the value function (Theorem 1), PI reaches the exact optimum in finitely many iterations. While VI requires $k=\log(100)/\log(1/\gamma)$ iterations to reduce error by a factor of 100 (Bertsekas, 1996), PI typically converges in 5–10 iterations regardless of $\gamma$. [^44] Ye (2011) proves PI is strongly polynomial with iteration count $O\!\left(\frac{|\mathcal{S}||\mathcal{A}|}{1-\gamma}\log\frac{|\mathcal{S}|}{1-\gamma}\right)$, resolving a long-standing conjecture. This bound is for fixed $\gamma$; Fearnley (2010) constructs examples requiring exponentially many iterations when $\gamma$ is allowed to vary with $|\mathcal{S}|$. [^45] For continuous-state problems discretized on a grid (the norm in economics), the finite-termination argument still applies to the discretized problem, but the number of grid points $n$ enters the bound. Santos and Rust (2004) establish a three-tier convergence result for PI applied to discretized dynamic programs: order $\approx 1.5$ globally for general interpolation schemes, quadratic convergence locally when the value function approximation is concave and piecewise linear, and superlinear convergence for general smooth interpolation. The formal error constants $C(h)$ in their quadratic bound degrade as the grid mesh $h\to 0$, but the iteration count is empirically independent of grid size. At $\gamma=0.90$, VI needs 44 iterations while PI needs only 5–8; at $\gamma=0.95$, VI requires 90 versus 5–8; at $\gamma=0.99$, VI needs 459 iterations while PI still converges in 5–10.
@@
Bertsekas (2022b) extends this interpretation to a broad class of dynamic programming problems. The Newton structure applies whenever the Bellman operator can be written as a pointwise maximum over linear operators: $T=\max_{\pi}T^{\pi}$. This includes not only infinite horizon problems with discounting but also optimal stopping problems (job search, option exercise), average cost optimization (inventory, queueing), and minimax formulations for adversarial settings. [^46] Rust (1996) surveys successive approximation and policy iteration methods for economic models, comparing their performance on the bus engine replacement problem. Zhang (2023) extends randomized policy iteration to multi-agent problems where the control is $m$-dimensional, reducing per-iteration complexity from exponential to linear in mm. [^47] These problems appear under different names such as “stochastic shortest path” for optimal stopping, “average-cost MDP” for long-run average optimization, and “model predictive control” (or receding-horizon control) for finite-horizon replanning. The practical implication is that algorithms with policy-improvement structure (evaluate a policy exactly or approximately, then improve) inherit Newton-like convergence behavior, while pure value-iteration methods (apply $T$ directly) are limited to linear convergence. [^48] Blackwell (1965) proves that for discounted MDPs with finite state and action spaces, a stationary deterministic policy $\pi:\mathcal{S}\to\mathcal{A}$ exists that is optimal for all initial states simultaneously. This “uniform optimality” has three implications: (1) the search space reduces from history-dependent or stochastic policies to static maps, justifying neural networks that condition only on current state; (2) the optimal policy is independent of the initial distribution $d_{0}$, so changing where episodes start does not require retraining; (3) PI and VI are guaranteed to converge to the same globally optimal policy regardless of initialization.